The time required for ten children to learn a particular motor skill was recorded as 9, 15, 23, 20, 16, 15, 24, 18, 10, and 20 minutes.Find the mean time to learn this task.
Answer #1

Solution:

Given that

The sample mean is

Mean   = (x / n)

= ( 9+15+23+20+16+15+24+18+10+20 / 10 )

= ( 170 / 10 )

= 17

Mean   = 17



=17010

=17